# HumanVid: Demystifying Training Data for Camera-controllable Human Image Animation

### v1

@@ Abstract
Human image animation involves generating videos from a character photo, allowing user control and unlocking potential for video and movie production. While recent approaches yield impressive results using high-quality training data, the inaccessibility of these datasets hampers fair and transparent benchmarking. Moreover, these approaches prioritize 2D human motion and overlook the significance of camera motions in videos, leading to limited control and unstable video generation.To demystify the training data, we present HumanVid, the first large-scale high-quality dataset tailored for human image animation, which combines crafted real-world and synthetic data. For the real-world data, we compile a vast collection of copyright-free real-world videos from the internet. Through a carefully designed rule-based filtering strategy, we ensure the inclusion of high-quality videos, resulting in a collection of 20K human-centric videos in 1080P resolution. Human and camera motion annotation is accomplished using a 2D pose estimator and a SLAM-based method. For the synthetic data, we gather 2,300 copyright-free 3D avatar assets to augment existing available 3D assets. Notably, we introduce a rule-based camera trajectory generation method, enabling the synthetic pipeline to incorporate diverse and precise camera motion annotation, which can rarely be found in real-world data. To verify the effectiveness of HumanVid, we establish a baseline model named CamAnimate, short for Camera-controllable Human Animation, that considers both human and camera motions as conditions. Through extensive experimentation, we demonstrate that such simple baseline training on our HumanVid achieves state-of-the-art performance in controlling both human pose and camera motions, setting a new benchmark. Code and data will be publicly available at https://github.com/zhenzhiwang/HumanVid/.
@@ 1 Introduction
Despite some promising progress [29, 86], human image animation presents two main challenges: the absence of a high-quality public dataset and the neglect of camera motions in human videos. Specifically, state-of-the-art approaches rely on private datasets for training similar models, underscoring the importance of datasets in this field. However, these datasets remain inaccessible, while accessible alternatives like TikTok [30] and UBC-Fashion [78] possess limitations in scale and quality, hindering fair and transparency evaluation and community development. Furthermore, despite leveraging enigmatic private datasets, these methods struggle to animate characters from new viewpoints with camera movement. They primarily rely on 2D pose extraction from static camera videos [72], neglecting the crucial aspect of camera motion. This design choice compromises video controllability and impedes high-quality character animation for complex motions.
@@
To begin with, we compile a vast collection of real-world videos from diverse scenes on copyright-free internet platforms. This uncurated dataset often contains noise, such as frequent lens switching, user-generated special visual effects, occlusions and object motions. To ensure high-quality videos, we employ a carefully designed rule-based filtering strategy to ensure that pixel motions in our collected videos are exclusively resulted from human and camera motions. Additionally, we utilize SLAM-based methods [66, 60] for accurate camera trajectory extraction and a precise pose estimator [72] to extract human pose sequences. This results in a remarkable collection of over 20K human-centric videos in 1080P resolution. Experimental results demonstrate that training models exclusively on this curated video dataset achieve state-of-the-art performances.
@@
Synthesizing diverse and high-fidelity human videos for animation is non-trivial. Existing 3D synthetic datasets for humans primarily focus on reconstruction or perception, resulting in limited appearance diversity, over-smoothed textures, and restricted camera motions [10, 73]. To overcome these limitations, we first augment our dataset with approximately 2,300 copyright-free 3D avatar assets. These assets undergo rigorous rigging using motions from motion-captured datasets [39] and open-source software [4], enabling a wide range of character shapes, appearances, and human motions. Notably, to compensate for the limited camera motions observed in real-world videos, we introduce an innovative rule-based camera trajectory system, enriching the diversity of camera movements in training data. Specifically, multiple camera locations are randomly sampled for the keyframes throughout the space, and each camera is purposefully directed toward the human face. These sampled key camera motions are then connected and smoothed, replicating common camera movements in videos and films. This design enables us to generate lifelike human videos with accurate annotations of human and camera motions. The illustration of our synthetic data is in Fig. 1. Our experiments conclusively demonstrate that utilizing our synthetic dataset significantly enhances animation, particularly regarding motion control.
@@
To validate the collected dataset, we incorporate camera control [23] into a widely-used model that only considers pose condition [29]. Our main contributions are as follows,
@@
  • To the best of our knowledge, we are the first to introduce a synthetic dataset combined with the first high-quality real-world video dataset for human image animation.
@@
  • We design a scalable pipeline that facilitates lifelike human video generation and provides accurate and diverse annotations of human and camera motions. This results in improved visual quality and enhanced control over human image animation.
@@
  • Through extensive experiments, we validate the effectiveness of each dataset component, establish a new state-of-the-art, and create a comprehensive, fair, and transparent evaluation benchmark for the field.
@@ 2 Related Work
Camera-aware Video Generation. As a significant component in video and movie production, camera viewpoint movement determines the content dynamics and the overall feeling of the audience. While many works focus on guiding video generative models with structural signals [18, 75, 63, 82, 32, 21], less attention has been paid to controlling the pose/viewpoint of camera in generating videos [70, 76]. To control camera motion with reference videos, MotionDirector [84] proposes a dual-path LoRA [28] adapter to decouple the motion and appearance learning and can roughly control camera movements to produce a surrounding shot. For more precise control, MotionCtrl [68] directly injects the camera extrinsic matrix to the temporal attention layer in pre-trained text-to-video models and can precisely specify the camera viewpoint by providing camera poses at inference. CameraCtrl [23] further enhances the controllability by representing the camera pose with Plücker ray embeddings [58, 36]. CamViG [40] explores the camera control in token-based video generator [35] by introducing camera embedding as a new modality.
@@ 3 Dataset
Figure 2: Flow chart of our synthetic data creation pipeline.
@@ 3.1 Synthetic Data Construction
The synthetic video data are rendered with one or more characters moving in various 3D scenes using diverse camera trajectories. Consequently, constructing the synthetic data involves three key steps: character creation, motion retargeting, and 3D scene and camera placement.
@@ 3.1.1 Character Creation
We create two types of characters for animating images in the diverse domains: (1) Human-like characters from SMPL-X [44] meshes and clothing. (2) Anime characters from user-uploaded assets. Diverse body shapes and skin textures, 3D clothing and textures are considered for highly varied human representation.
Figure 4: Illustration of the SMPL-X characters rendered with background.
Body shapes and skin tone. For human-like characters, we sample body shapes from a diverse set of 271 body shapes with different BMI collected from the ARGOA [43] and CAESAR [49] datasets following Bedlam [10]. To reduce the gender and ethnicity bias, we use 50 female and 50 male commercial skin albedo textures collected from Meshcapade [19] with a resolution of 4096 $\times$ 4096, spanning over seven ethnic groups.
3D Clothing and textures. To generate realistic human videos, it’s crucial to have 3D clothing motions that are physically plausible and consistent with human body movements. For instance, the LSMPL-X representation from Synbody [73] adds a clothing layer to SMPL-X [44], but lacks realistic physics simulation for clothing motion, leading to unnatural movements in loose-fitting clothes like dresses. We collect 111 unique outfits, including T-shirts, sweaters, coats, jeans and skirts from Bedlam [10] dataset, and use commercial simulation software [1] to obtain realistic clothing deformations. On top of realistic meshes of human mesh and clothing from physics simulation, 1691 unique clothing textures are used for diverse clothing appearances. We shows examples of clothed SMPL-X characters in Fig. 4.
Anime characters. To enhance the diversity of characters in our synthetic data, we focused on VRoidHub [6], a platform designed for sharing and showcasing 3D character models. Creators on VRoidHub have uploaded a vast array of intricate 3D character models, featuring diverse appearances, clothing styles, and hairstyles. From this rich repository, we manually selected 2,387 characters for further rendering.
@@ 3.1.2 Motion Retargeting
Anime characters. Conversely, anime character assets can have diverse skeleton lengths. We utilize the automatic re-targeting software Rokoko to transfer existing motions to the anime character assets. The clothing and hair are treated as part of the body, so their motion is also determined by source motions. The illustration of transferring motion to anime assets is shown in Fig. 3.
@@
3D Scenes. The realistic and diverse 3D scene backgrounds for synthetic video are constructed from about 100 panoramic HDRI images [2] or high-quality 3D scenes to cover both indoor and outdoor environment. We manually select panorama images with flat ground for characters to move on, while avoiding excessive scene components that might lead to unnatural human-scene interactions. We also exclude images with uniform visual patterns across different views, such as grasslands, deserts, or farms. The selected panorama backgrounds feature high-quality, complex texture details that highlight varying background textures from different camera angles. We provide more details in the supplementary materials.
Camera Trajectory Design. Unlike [10, 73, 43], our dataset highlights rich and diverse camera trajectories in human-centric videos. Each camera trajectory consists of a sequence of $6$-DoF translations and rotations. We carefully design a rule-based camera motion generation pipeline to obtain diverse trajectories. This pipeline randomly sample camera locations adaptive to human positions and orientations in the keyframes, and use spline interpolation to get smooth camera locations and rotations in the whole sequence. Specifically, in each keyframe scene space, we randomly sample camera locations within a semi-cylinder of radius $\in[3m,5m]$ and height $\in[0.6m,1.2m]$ in front of the human. Then, we set the camera orientation’s yaw and pitch to point at the person. To create a more natural camera trajectory that smoothly follows the person, we adjust the camera’s position by adding the human’s position offset from the keyframe to the camera position in each frame. Finally, we also sample the roll of camera rotation $\in[-30^{\circ},30^{\circ}]$ in keyframes. Our design of camera keyframe sampling enables all types of camera trajectories, significantly enhancing the camera trajectory diversity and cinematic effect of human videos compared to existing video datasets.
Rendering and Annotations. We render the image sequences of SMPL-X characters using UE5 game engine and the built-in movie render function (Movie Render Queue) for high-quality images. The anime characters are rigged and rendered with blender. With our synthetic data source, a variety of ground-truth annotations, including camera trajectories, human skeletons, segmentation masks, depth maps and normal maps, could be obtained without manual efforts. Although our dataset is curated for human video generation, these ground-truth annotations could also be useful for other downstream applications.
@@ 3.2 Internet Data Curation
To enhance the appearance diversity on top of synthetic videos, we collect real human-centric videos from copyright-free internet platforms [5] with the same distribution of synthetic data: the pixel motions in such videos is only resulted from human skeleton motion or camera motion, not object movements or background dynamics.
We utilize the Pexels API [5] to scrape data based on around 100 keywords and employed a pose detector [72] to analyze the data. The pose detector focused on measuring the upper body keypoints’ confidence, the ratio of the largest human bounding box over the frame $r$, the average number of humans present in each frame $n$, and the average motion (position offsets) of the keypoints $\Delta\bar{\mathbf{p}}$. With these statistics, we apply a specific filtering criteria: a) the human should occupy the main part of the image ($r>0.07$); b) there should be few people ($n\leq 4$); c) there should be a noticeable motion in keypoints to remove static videos ($\Delta\bar{\mathbf{p}}>0.01$); d) No exits, entrances or occlusions of individuals in videos ($n\in\mathbb{Z}$). As a result, we collect around 20k high-quality, real human-centric video clips with various appearances.
Camera Trajectory Estimation. Reconstructing global camera trajectories from in-the-wild videos is a challenging problem. Following the curation of human-centric videos from previous steps, we adopt TRAM [66] to utilize a SLAM method [60] for recovering camera extrinsic parameters from in-the-wild videos with explicit human movement. To ensure camera parameters are robust to dynamic humans, we employ a masking technique [34] that removes dynamic regions from both input images and dense bundle adjustment steps. By compelling SLAM to rely solely on the background for camera estimation from the outset, we mitigate the risk of catastrophic failure. To convert camera estimation to metric scale, we leverage semantic cues from the background by utilizing noisy depth predictions [9]. Consequently, we recover accurate, metric-scale camera motion that serves as an optimal camera condition for training diffusion models. For videos with texture-less backgrounds, where the SLAM system reports large reconstruction errors, we empirically set such samples to static cameras. Additionally, we filter out videos with very large rotations or translations, such as cycling, or those with sudden shot changes, as these videos fall outside the scope of human image animation.
@@
Ours Synthetic (SMPL-X) 50k 8M 1080P Ground Truth Ground Truth
Ours Synthetic (Anime) 50k 4M 1080P Ground Truth Ground Truth
Statistics. As shown in Tab. 1, our Real Internet dataset, with over 20k clips and 10M frames at 1080P resolution, significantly surpasses existing datasets like TikTok [30], UBC-Fashion [78] and IDEA-400 [37] in both size and resolution. Additionally, our synthetic dataset from SMPL-X and Anime characters, is 5×5$\times$5 × larger in scale compared to Bedlam [10], providing accurate camera and human pose groundtruth, and more diverse camera trajectories.
Figure 5: Illustration of archtecture of our CamAnimate.

### v3

@@ Abstract
Human image animation involves generating videos from a character photo, allowing user control and unlocking the potential for video and movie production. While recent approaches yield impressive results using high-quality training data, the inaccessibility of these datasets hampers fair and transparent benchmarking. Moreover, these approaches prioritize 2D human motion and overlook the significance of camera motions in videos, leading to limited control and unstable video generation. To demystify the training data, we present HumanVid, the first large-scale high-quality dataset tailored for human image animation, which combines crafted real-world and synthetic data. For the real-world data, we compile a vast collection of real-world videos from the internet. We developed and applied careful filtering rules to ensure video quality, resulting in a curated collection of 20K high-resolution (1080P) human-centric videos. Human and camera motion annotation is accomplished using a 2D pose estimator and a SLAM-based method. To expand our synthetic dataset, we collected 10K 3D avatar assets and leveraged existing assets of body shapes, skin textures and clothings. Notably, we introduce a rule-based camera trajectory generation method, enabling the synthetic pipeline to incorporate diverse and precise camera motion annotation, which can rarely be found in real-world data. To verify the effectiveness of HumanVid, we establish a baseline model named CamAnimate, short for Camera-controllable Human Animation, that considers both human and camera motions as conditions. Through extensive experimentation, we demonstrate that such simple baseline training on our HumanVid achieves state-of-the-art performance in controlling both human pose and camera motions, setting a new benchmark. Demo, data and code could be found in the project website: https://humanvid.github.io/.
@@ 1 Introduction
Despite recent advances [28, 88], human image animation presents two main challenges: the absence of a high-quality public dataset and the neglect of camera motions in human videos. Specifically, state-of-the-art approaches rely on private datasets for training similar models, underscoring the importance of datasets in this field. However, these datasets remain inaccessible, while accessible alternatives like TikTok [29] and UBC-Fashion [80] possess limitations in scale and quality, hindering fair and transparency evaluation and community development. Furthermore, despite leveraging private datasets, these methods struggle to animate characters from new viewpoints with camera movement. They primarily rely on 2D pose extraction from static camera videos [75], neglecting the crucial aspect of camera motion. This design choice compromises video controllability and impedes high-quality character animation for complex motions.
@@
To begin with, we compile a vast collection of real-world videos from diverse scenes on copyright-free internet platforms. This uncurated dataset often contains noise, such as frequent shot changes, user-generated special visual effects, occlusions and object motions. To ensure high-quality videos, we employ a carefully designed rule-based filtering strategy to ensure that pixel motions in our collected videos are exclusively resulted from human and camera motions. Additionally, we utilize SLAM-based methods [69, 62] for accurate camera trajectory extraction and a precise pose estimator [75] to extract human pose sequences. This results in a remarkable collection of over 20K human-centric videos in 1080P resolution. Experimental results demonstrate that training models exclusively on this curated video dataset achieve state-of-the-art performances.
@@
Synthesizing diverse and high-fidelity human videos for animation is non-trivial. Existing 3D synthetic datasets for humans primarily focus on reconstruction or perception, resulting in limited appearance diversity, over-smoothed textures, and restricted camera motions [10, 76]. To overcome these limitations, we first augment our dataset with approximately 10K copyright-free 3D avatar assets. These assets undergo rigorous rigging using motions from motion-captured datasets [40] and open-source software [4], enabling a wide range of character shapes, appearances, and human motions. Notably, to compensate for the limited camera motions observed in real-world videos, we introduce an innovative rule-based camera trajectory system, enriching the diversity of camera movements in training data. Specifically, multiple camera locations are randomly sampled for the keyframes throughout the space, and each camera is purposefully directed toward the human face. We connect and smooth these sampled camera keyframes to create natural camera movements similar to those found in professional videos and films. This design enables us to generate lifelike human videos with accurate annotations of human and camera motions. The illustration of our synthetic data in the rendering process is shown in Fig. 1. Our experiments conclusively demonstrate that utilizing our synthetic dataset significantly enhances animation, particularly regarding motion control.
@@
To validate the collected dataset, we incorporate camera control [22] into a widely-used human video generation model that only considers pose condition [28]. Our main contributions are as follows,
@@
  • To the best of our knowledge, HumanVid is the first large-scale video dataset for human image animation. It contains both high-quality Internet videos with diverse appearance and synthetic videos with accurate human pose and camera pose annotations.
@@
  • We design a scalable rendering pipeline from Unreal Engine 5 that facilitates lifelike human video generation and provides accurate annotations of human and camera motions.
@@
  • Through extensive experiments, we validate the effectiveness of each dataset component, establish a new state-of-the-art, and create a comprehensive and transparent evaluation benchmark for the field.
@@ 2 Related Work
Camera-aware Video Generation. As a significant component in video and movie production, camera viewpoint movement determines the content dynamics and the overall feeling of the audience. While many works focus on guiding video generative models with structural signals [17, 77, 66, 83, 32, 20], less attention has been paid to controlling the pose/viewpoint of camera in generating videos [73, 78]. To control camera motion with reference videos, MotionDirector [85] proposes a dual-path LoRA [27] adapter to decouple the motion and appearance learning and can roughly control camera movements to produce a surrounding shot. For more precise control, MotionCtrl [71] directly injects the camera extrinsic matrix to the temporal attention layer in pre-trained text-to-video models and can precisely specify the camera viewpoint by providing camera poses at inference. CameraCtrl [22] further enhances the controllability by representing the camera pose with Plücker ray embeddings [60, 36]. CamViG [41] explores the camera control in token-based video generator [34] by introducing camera embedding as a new modality. Furthermore, JAWS [65] and Jiang et. al. [30] explore video generation via cinematic transfer from existing videos. To the best of our knowledge, our paper is the first to introduce camera control into the human-centric video generation task.
@@ 3.1 Synthetic Data Construction
The synthetic video data are rendered with one character moving in various 3D scenes using diverse camera trajectories. Consequently, constructing the synthetic data involves three key steps: character creation, motion retargeting, and 3D scene and camera placement.
@@ 3.1.1 Character Creation
We create two types of characters in the diverse domains: (1) Human-like characters from SMPL-X [45] meshes and clothing. (2) Anime characters from user-uploaded assets. Diverse body shapes and skin textures, 3D clothing and textures are considered for highly varied human representation.
Figure 2: Illustration of the clothed SMPL-X characters rendered with diverse backgrounds.
Body shapes and skin tone. For human-like characters, we sample body shapes from a diverse set of 271 body shapes with different BMI collected from the ARGOA [44] and CAESAR [50] datasets following Bedlam [10]. To reduce the gender and ethnicity bias, we use 50 female and 50 male commercial high-resolution skin albedo textures from Meshcapade [18] with seven ethnic groups.
3D Clothing and textures. To generate realistic human videos, it’s crucial to have 3D clothing motions that are physically plausible and consistent with human body movements. For instance, the LSMPL-X representation from Synbody [76] adds a clothing layer to SMPL-X [45], but lacks realistic physics simulation for clothing motion, leading to unnatural movements in loose-fitting clothes like dresses. We collect 111 unique outfits, including T-shirts, sweaters, coats, jeans and skirts from Bedlam [10] dataset, and use commercial simulation software [1] to obtain realistic clothing deformations. On top of realistic meshes of human mesh and clothing from physics simulation, 1,691 unique clothing textures are used for diverse clothing appearances. We shows examples of clothed SMPL-X characters in Fig. 2.
Anime characters. To enhance the diversity of characters in our synthetic data, we focused on VRoidHub [6], a platform designed for sharing 3D character models. Creators on VRoidHub have uploaded a vast array of intricate 3D character models, featuring diverse appearances, clothing styles, and hairstyles. From this rich repository, we manually selected 10K characters for rendering.
@@ 3.1.2 Motion Retargeting
Anime characters. Conversely, anime character assets could have diverse skeleton lengths. We utilize the re-targeting software [4] to transfer existing motions to the anime character assets. The clothing and hair are treated as part of the body, so their motion is also determined by source motions.
Figure 3: Illustration of the light condition control in the rendering process.
@@
3D Scenes. The realistic and diverse 3D scene backgrounds for synthetic video are constructed from about 100 panoramic HDRI images [2] or high-quality 3D scenes to cover both indoor and outdoor environment. We manually select panorama images with flat ground for characters to move on, while avoiding excessive scene components that might lead to unnatural human-scene interactions. We also exclude images with uniform visual patterns across different views, such as grasslands, deserts, or farms. The selected panorama backgrounds feature high-quality, complex texture details that highlight varying background textures from different camera angles.
Camera Trajectory Design. Unlike [10, 76, 44], our dataset highlights rich and diverse camera trajectories in human-centric videos. Each camera trajectory consists of a sequence of $6$-DoF translations and rotations. We carefully design a rule-based camera motion generation pipeline to obtain diverse trajectories. This pipeline randomly sample camera locations adaptive to human positions and orientations in the keyframes, and use spline interpolation to get smooth camera locations and rotations in the whole sequence. Specifically, in each keyframe, we randomly sample camera locations within a semi-cylinder of radius $\in[3m,5m]$ and height $\in[0.6m,1.2m]$ in front of the human. Then, we set the camera orientation’s yaw and pitch to point at the person. To create a more natural camera trajectory that smoothly follows the person, we adjust the camera’s position by adding the human’s position offset from the keyframe to the camera position in each frame. Finally, we also sample the roll of camera rotation $\in[-30^{\circ},30^{\circ}]$ in keyframes. Our design of camera keyframe sampling enables all types of camera trajectories, significantly enhancing the camera trajectory diversity and cinematic effect of human videos compared to existing video datasets.
Rendering and Annotations. We render the image sequences of SMPL-X characters using UE5 game engine and the built-in movie render function (Movie Render Queue) for high-quality images. The anime characters are rigged and rendered with blender. With our synthetic data source, a variety of ground-truth annotations, including camera trajectories, human skeletons, segmentation masks, depth maps and normal maps, could be obtained without manual efforts. Although our dataset is curated for human video generation, these ground-truth annotations could also be useful for other downstream applications. Thanks to the Unreal Engine, we could render photo-realistic videos with controllable human appearance and motion, scene, and light condition. As shown in Fig. 3, the light direction in the rendering process could be controlled from $-17^{\circ}$ to $13^{\circ}$.
@@ 3.2 Internet Data Curation
To enhance the appearance diversity of synthetic videos, we collect real human-centric videos from copyright-free internet platforms [5], where the pixel motions in videos is only resulted from human skeleton motion or camera motion, without any object movements or background dynamics. We utilize the Pexels API [5] to scrape data based on around 100 keywords and employed a pose detector [75] to analyze the data. The pose detector focused on measuring the upper body keypoints’ confidence, the ratio of the largest human bounding box over the frame $r$, the average number of humans present in each frame $n$, and the average motion (position offsets) of the keypoints $\Delta\bar{\mathbf{p}}$. With these statistics, we apply a specific filtering criteria: a) the human should occupy the main part of the image ($r>0.07$); b) there should be few people ($n\leq 4$); c) there should be a noticeable motion in keypoints to remove static videos ($\Delta\bar{\mathbf{p}}>0.01$); d) No exits, entrances or occlusions of individuals in videos ($\mid n-\mathrm{round}(n)\mid<0.01$). As a result, we collect around 20k high-quality, real human-centric video clips with various human and scene appearances.
Camera Trajectory Estimation. Reconstructing global camera trajectories from in-the-wild videos is a challenging problem. For the curated human-centric videos, we adopt TRAM [69] to utilize a SLAM method [62] for recovering camera extrinsic parameters from in-the-wild videos with explicit human movement. To ensure camera parameters are robust to dynamic humans, we employ a masking technique [33] that removes dynamic regions from both input images and dense bundle adjustment steps. To prevent major estimation errors, we configure the SLAM system to use only background features for camera motion estimation. To convert camera estimation to metric scale, we leverage semantic cues from the background by utilizing noisy depth predictions [9]. Consequently, we recover accurate, metric-scale camera motion that serves as an optimal camera condition for training diffusion models. When videos have backgrounds lacking texture and the SLAM system cannot accurately reconstruct camera motion, we treat these cases as having static cameras. Additionally, we filter out videos with very large rotations or translations, such as cycling, or those with sudden shot changes, as these videos fall outside the scope of human image animation.
@@
Ours Synthetic (SMPL-X) 50k 8M 720P Ground Truth Ground Truth
Ours Synthetic (Anime) 25k 2M 1080P Ground Truth Ground Truth
Table 2: Statistics of the diversity of appearance, motion and scene in HumanVid.
Dataset Split #Subject #Motion #Scene Avg. Clip Length
Internet videos 24,012 24,012 19,688 (= #video) 16.65s
Synthetic (SMPL-X)
271 (body shapes)
$\times$ 100 (skin textures)
$\times$ 1,691 (clothings)
2,311
100 (HDRIs)
+ 587 (3D scenes)
6.34s
Synthetic (Anime) 10K (anime assets) 40
100 (HDRIs)
+ 93 (3D scenes)
3.2s
Statistics. As shown in Tab. 1, our Real Internet dataset, with over 20k clips and 10M frames at 1080P resolution, significantly surpasses existing datasets like TikTok [29], UBC-Fashion [80] and IDEA-400 [37] in both size and resolution. Additionally, our synthetic dataset from SMPL-X and Anime characters, is 5×5$\times$5 × larger in scale compared to Bedlam [10], providing accurate camera and human pose groundtruth, and more diverse camera trajectories. For statistics of our HumanVid, Tab. 2 shows the statistics of human appearance, motion and scene.
Figure 4: Illustration of the network architecture of our proposed CamAnimate.